OpenStudy (anonymous):

Prove that for all integers a, a^2 >=a

OpenStudy (anonymous):

what method are you supposed to use?

OpenStudy (anonymous):

it is certainly true for all negative integers, because \(a^2>0\)

OpenStudy (anonymous):

right well its certainly true no matter what but proving it is the tough part for me

OpenStudy (anonymous):

i'm pretty sure i'm supposed to split it into 3 cases: a is a natural number, -a is a natural numer, and a=0

OpenStudy (anonymous):

and a is a natural number is the one i'm stuck on

OpenStudy (anonymous):

for positive integers you could use induction

OpenStudy (anonymous):

i've heard that before, but i can't remember what it is

OpenStudy (anonymous):

try this since \(n(n-1)>0\) if both \(n\) and \(n-1\) are, then \[n^2-n>0\] and so \(n^2>n\) handle the case 0 and one separately

OpenStudy (anonymous):

so this means n>2 though in this example right?

OpenStudy (anonymous):

i mean 1(1-1) isn't greater than 0

OpenStudy (anonymous):

no problem with 1 since \(1^2=1\)

OpenStudy (anonymous):

but i'm just confused as to how you can assume n(n-1)>0

OpenStudy (anonymous):

i guess just because you're assuming n>1?

OpenStudy (tkhunny):

You do not need to assume that n > 1. For n < 0, it is relatively obvious n < 0 and \(n^{2} > 0\) For n = 0 or 1, we have \(n^{2} = n\) The only hard part is for n > 1 There are a few ways to proceed. I like this one. \(4 = 2^{2} > 2\) If \(a^{2} > a\), we have If \(a^{2} + 1 > a + 1\), and then If \((a+1)^{2} = a^{2} + 2a + 1> a^{2} + 1 > (a + 1)\), and the proof is complete.

OpenStudy (anonymous):

that bottom line is a little fuzzy, what are you doing there?

OpenStudy (anonymous):

like where's the (a+1)^2 coming from?

OpenStudy (tkhunny):

The bottom line is the meat. It came from factoring. Algebra 1. Also, @Satellite73 is right on target. After demonstrating for n < 2, just like I did, n*(n-1) > 0. It is simple and clear.

OpenStudy (anonymous):

I know, i understood what he said, and i understand that you factored to get from the first to second step in the bottom row, but then the 2a just disappears...

OpenStudy (tkhunny):

I didn't get from the first step to the second step. You're reading it backwards. One piece at a time. \(a^{2} > a\) \(a^{2} + 1 > a + 1\) -- Just addition \(a^{2} + 2a + 1 > a^{2} + 1 > a + 1\) -- Just addition \((a+1)^{2} = a^{2} + 2a + 1 > a^{2} + 1 > (a + 1)\) -- Factoring It's not magic. I'm just heeding the inequality. Increasing one side by a positive value maintains the inequality. Factoring an expression - not changing the value at all - maintains the inequality.

OpenStudy (anonymous):

I see what you did, I didn't realize you were adding a third part to the inequality, thats a really cool way of doing it. and you can do that since you're assuming a>1, right?

OpenStudy (tkhunny):

Correct. We needed to get rid of the negative numbers (since adding 2a might violate the inequality) and the equalities (because they are annoying). It's called "Mathematical Induction". We used this: If \(a^{2} > a\), then \((a+1)^{2} > (a+1)\) and proved it by direct construction. Since 'a' is arbitrary, we just proved strick inequality for all Natural Numbers > 1.

OpenStudy (anonymous):

Ok cool, thanks so much. that makes a lot of sense now.

OpenStudy (anonymous):

And thank you too satellite, you helped a lot as well!